OpenStudy (lizz123):

Is this correct y=-1x+1 (5,0)(2,-3) Write an equation in slope intercept form the line passing through each pair of points. @iGreen.

jimthompson5910 (jim_thompson5910):

Let's check the answer

jimthompson5910 (jim_thompson5910):

if y = -1x + 1 was the answer, then both points must be on the line.

OpenStudy (alexandervonhumboldt2):

uhh no not correct

OpenStudy (alexandervonhumboldt2):

the slope is 1 not -1

jimthompson5910 (jim_thompson5910):

let's see if (5,0) is on the line y = -1x+1 0 = -1(5) + 1 ... plug in (x,y) = (5,0) 0 = -5 + 1 0 = -4 ... this is false, so (5,0) is NOT on the line y = -1x+1

OpenStudy (anonymous):

very sorry incorrect

OpenStudy (lizz123):

thanks @jim_thompson5910 I have couple more. Can I tag you?

jimthompson5910 (jim_thompson5910):

sure